XYZ Co. operates in a competitive market. Its Total Product (Q) is given as Q=f(L)= 3L, and it



takes the wage and price as given. Derive the firm's short-run demand for labor as a function of w



and p. How much labor will the firm hire if W=₵25 and P=₵150?

1
Expert's answer
2021-11-04T13:49:18-0400

The short run demand curve is obtained by dividing the nominal demand curve by the product price:

"MP_L=\frac{W}{p}=w"

"MP_L=3L"

"P=150"

"W=25"

"MP_L\times p=w"

"3L\times150=25"

"L=0.056"

"Q=f(L) =3L"

"\implies Q=3(0.056)=0.168"

0.168 represents the quantity of labor that the firm will hire.